# Ensemble noise properties of the European Pulsar Timing Array

### v2

@@ 3.3 Distribution of noise parameters as a mixture model
One may also be interested whether the distribution of pulsar noise parameters is more complex. For example, in Figure 1 in Goncharov et al. (2021a), one may notice the clustering of noise parameters in two areas. First, around $\gamma\approx 1$. Second, a cluster of pulsars around $\gamma=13/3$ of the gravitational wave background. A similar clustering is observed in the PTA noise analysis by the NANOGrav Collaboration (2023b), where it is explained that pulsars with $\gamma\approx 13/3$ contribute to the common-spectrum process in full-PTA analysis. An analogous Figure is shown in Reardon et al. (2023), although the presence of two over-densities is not obvious there. Although we have taken steps to marginalize the contribution of the common-spectrum process and uncover properties of pulsar-intrinsic noise, it is nevertheless useful to consider the possibility of a mixture model of a Gaussian distribution $\mathcal{N}$ and a Uniform distribution $\mathcal{U}$, where a broader Uniform distribution may fit potential outliers. The density of such a distribution is then $\nu\mathcal{N}+(1-\nu)\mathcal{U}$, where $\nu\in[0,1]$ is the contribution of the normal distribution.
@@ 3.4 Summary of the results
Some ensemble properties of spin noise (SN) are similar to DM noise, whereas some properties are different. For spin noise, the normal distribution fits the data as well as the uniform distribution. However, the covariance between $\lg A_{\text{SN}}$ and $\gamma_{\text{SN}}$ plays an important role. In Figures 1(a) and 2(a), one may notice a diagonal trend which results in finding a covariance parameter $\rho_{\text{SN}}\approx-1$. The model with a lack of covariance, $\rho_{\text{SN}}=0$ is disfavored with $\ln\mathcal{B}=6$. The covariance between $\lg A_{\text{SN}}$ and $\gamma_{\text{SN}}$ can be explained by noticing that it follows a line of equal noise power. In contrast, for DM noise, we do not find strong evidence for covariance between $\lg A_{\text{DM}}$ and $\gamma_{\text{DM}}$. With a lack of covariance, the uniform model is strongly preferred over the Gaussian model with a log Bayes factor of $6$. The mixture model is disfavored by our observations of both spin noise and DM noise, indicating the lack of outlying noise terms.
Table 2: Results of hyperparameter estimation and hierarchical model selection for ensemble pulsar noise properties of the European Pulsar Timing Array. DM noise and spin noise are parametrized by $\bm{\theta}=(\lg A,\gamma)$. We present the results for the following models of pulsar noise parameter distributions: a uniform distribution $\mathcal{U}$, a normal distribution $\mathcal{N}$ with and without covariance between $\lg A$ and $\gamma$, a mixture model of a normal and a uniform distribution. Columns $\mathcal{P}(\bm{\Lambda}|\bm{\delta t})$ contain maximum-aposteriori hyperparameter values with $1\sigma$ credible levels. Columns $\ln\mathcal{B}$ contain log Bayes factors in favour of a model of interest against the uniform distribution model. The fact that they are negative means that the uniform model is the best fit.
@@
$\sigma_{\gamma_{\text{SN}}}$ $1.48^{+0.00}_{-0.00}$ $\sigma_{\gamma_{\text{DM}}}$ $0.97^{+0.40}_{-0.24}$
@@ 4 Caveats of the circular analysis
Circular analysis or “double dipping” refers to the usage of data to inform on the model which is then applied to analyse the same data. Such a practice may lead to the underestimation of measurement uncertainty and systematic errors. This point is stressed by van Haasteren (2024, in prep.). The author argues against the approach of determining which pulsar spin noise terms to include in the full-PTA analysis based on (a lack of) evidence for these terms in single-pulsar noise model selection. Indeed, double dipping is not a proper statistical treatment, it should be abandoned in favour of the proper model selection and model averaging performed in one global fit, as recommended by the author. However, to give a fair overview of the problem, we also show that circular analysis does not always yield incorrect results in analyses of PTA data.
Before we discuss the application of hierarchical inference in the subsections below, we would like to point out that it is not clear from van Haasteren (2024, in prep.) that the aforementioned approach of selecting spin noise terms for subsequent full-PTA analyses has led to underestimation of measurement uncertainties reported by PTAs. For model averaging, the prior odds between the existence or absence of common red noise are proposed to be 50%-50% in van Haasteren (2024, in prep.). For the case of including a red noise term to every pulsar, the prior odds will be arbitrary and vary from pulsar to pulsar which is clear from the equations in the Appendix in van Haasteren (2024, in prep.). However, data informs on these odds, so they should become a model (hyper-)parameter to avoid prior misspecification, the same problem discussed in this work. These parametrised odds will act as model selection. In the limit where there is no measurable red noise in pulsars and the data informs on it sufficiently well, this will lead to the elimination of the contribution of one of the models. Because different noise models lead to different measurement uncertainties for $(\lg A_{\text{CP}},\gamma_{\text{CP}})$, it is possible that the uncertainty introduced by enforcing red noise to every pulsar is suboptimal.
@@ 4.1 A comparison between circular analysis, incorrect priors, and a proper analysis
One example of circular analysis in the context of our hierarchical inference is finding best-fit $\bm{\Lambda}$ and using it for the gravitational wave search. It also leads to resolving prior misspecification but potentially at the cost of a reduced measurement uncertainty for $(\lg A,\gamma)$ of the gravitational wave background. The reduction may come thanks to simply ignoring a part of the intrinsic measurement uncertainty for $\bm{\Lambda}$, so a regular noise fluctuation is more likely to render our estimate of $(\lg A,\gamma)$ to be inconsistent with the true value. In Figure 3, we show three measurements of $(\lg A,\gamma)$ of the common-spectrum process in the EPTA data. Blue dashed contours correspond to a posterior obtained with the standard uniform spin noise priors, $\pi(\bm{\theta}|\bm{\Lambda})=\mathcal{U}(\bm{\theta}|\varnothing)$. Red dotted contours correspond to $\pi(\bm{\theta}|\bm{\Lambda})=\mathcal{N}(\bm{\theta}|\bm{\mu}_{0},\bm{\sigma}_{0},\rho=0)$, where $(\bm{\mu}_{0},\bm{\sigma}_{0})$ are the values obtained for this model from Table 2. Therefore, the red dotted line corresponds to a circular analysis. Filled green contours correspond to a proper (global-fit) measurement obtained based on Equation 9, it is discussed in more detail in the companion paper by Goncharov et al. (2024, in prep.). The measurement uncertainty of $(\lg A,\gamma)$ obtained with circular analysis matches that obtained with a full hierarchical analysis. It is even $5$-$8$% larger, suggesting that in the current setting circular analysis overestimates the intrinsic measurement uncertainty. Note, however, that the circular analysis shifts the maximum-aposteriori value of $(\lg A,\gamma)$, so the use of it is still not recommended.
@@
Figure 3: A comparison between circular analysis (red dotted contours), incorrect broad uniform priors (blue dashed contours), and a proper hierarchical analysis (green contours) of the common-spectrum process associated with the gravitational wave background.
4.2 A toy example: circular analysis of common red noise
@@
Keeping in mind the caveats above, for the test purposes we also perform the estimation of hyperparameters $\bm{\Lambda}$ for the mixture model of two Gaussian distributions where hyper-priors on the second component of the mixture correspond to EPTA measurement of the common-spectrum process parameters EPTA Collaboration & InPTA Collaboration (2023b). This exercise remains useful to explore the interplay between pulsar-intrinsic noise and the common-spectrum process and to understand the performance of the model. For the two-Gaussian mixture model, $\pi(\lg A_{\text{SN}},\gamma_{\text{SN}}|\bm{\Lambda}_{2\mathcal{N}})$, and $\bm{\Lambda}_{2\mathcal{N}}=(\mu_{\lg A_{\text{SN}}}^{1,2}$, $\sigma_{\lg A_{\text{SN}}}^{1,2}$, $\mu_{\gamma_{\text{SN}}}^{1,2}$, $\sigma_{\gamma_{\text{SN}}}^{1,2}$, $\rho_{\text{SN}}^{1,2}$, and $\nu_{\mathcal{N}})$. Indices $(1,2)$ refer to each of the two components in a mixture, while $\nu$ corresponds to a fraction of the first mixture component in the total prior probability. For the second component to correspond to the common-spectrum process in the EPTA data, we impose a Gaussian hyperprior determined by the measurement uncertainty on $(\lg A,\gamma)$ reported by the EPTA for our data. We find that the fraction of the second normal component is mostly consistent with one and inconsistent with zero. However, hyperparameters of the second mixture component – mean and (co-)variance – are well-constrained. Interestingly, we also find a marginal excess posterior density at $(\sigma_{\lg A_{\text{SN}}}^{2},\sigma_{\gamma_{\text{SN}}}^{2})=(0,0)$. This result is related to Goncharov et al. (2022)’s measurement of $\sigma_{\lg A}$ to be consistent with zero. When removing the common-spectrum term with $\gamma=13/3$ from our pulsar-intrinsic noise model, the posterior density increases, as expected. This is shown in Figure 4.
@@
Figure 4: Modelling of the common-spectrum process as a narrow normally-distributed cluster of pulsar noise using data-informed hyperpriors $\pi(\bm{\Lambda})$. We assume that pulsar noise parameters $\lg A$ and $\gamma$ are distributed as a mixture model of two Gaussians, where hyperparameters of the second Gaussian are chosen to match the measurement uncertainty on the common spectrum process reported by the EPTA Collaboration & InPTA Collaboration (2023b). Parameters $\sigma$ with respective indices correspond to standard deviations of the distribution that governs $\lg A$ and $\gamma$. Consistency of $\sigma$ with zero has first been shown in Goncharov et al. (2022) using a more rigorous although less flexible methodology. Parameter $\nu_{\mathcal{N}}$ corresponds to the fraction of the first, wider normal distribution in a model. Solid lines correspond to including the second $\gamma=13/3$ red noise term to pulsars to account for the presence of the common-spectrum stochastic process, dashed lines correspond to not including it. As expected, the contribution of the common-spectrum process becomes more visible in the latter case at $\sigma_{\lg A,\gamma}=0$ and $\nu_{\mathcal{N}}\approx 0.3$.
@@ 5 Conclusions
We performed inference of ensemble properties of spin noise and DM noise in the 25-year version of the second data release of the European Pulsar Timing Array (EPTA). Overall, we find that the standard uniform priors used in the previous analysis are too wide and not representative of the observations. However, a parametrised uniform prior distribution is a good fit. We, therefore, recommend correctly accounting for ensemble pulsar noise properties using Equation 9 in future full-PTA analyses to avoid systematic errors (parameter estimation and model selection biases).
@@ Acknowledgements
The data used for our study, EPTA DR2 by the (EPTA Collaboration, 2023), is available at zenodo. The code to reproduce the results of this analysis and the analysis by Goncharov et al. (2022) is available in the epta branch at github.com/bvgoncharov/pta_gwb_priors. Other data may be provided by the corresponding author upon request. The PTA likelihood is incorporated in enterprise (Ellis et al., 2020) and posterior sampling is performed using ptmcmcsampler (Ellis & van Haasteren, 2017) and dynesty (Speagle, 2020).
Appendix A Numerical marginalisation over hyperparameters

### v3

@@ 3.3 Distribution of noise parameters as a mixture model
One may also be interested whether the distribution of pulsar noise parameters is more complex. For example, in Figure 1 in Goncharov et al. (2021a), one may notice the clustering of noise parameters in two areas. First, around $\gamma\approx 1$. Second, a cluster of pulsars around $\gamma=13/3$ of the gravitational wave background. A similar clustering is observed in the PTA noise analysis by the NANOGrav Collaboration (2023b), where it is explained that pulsars with $\gamma\approx 13/3$ contribute to CP in full-PTA analysis. An analogous Figure is shown in Reardon et al. (2023), although the presence of two over-densities is not obvious there. Although we have taken steps to marginalize the contribution of the CP and uncover properties of pulsar-intrinsic noise, it is nevertheless useful to consider the possibility of a mixture model of a Gaussian distribution $\mathcal{N}$ and a Uniform distribution $\mathcal{U}$, where a broader Uniform distribution may fit potential outliers. The density of such a distribution is then $\nu\mathcal{N}+(1-\nu)\mathcal{U}$, where $\nu\in[0,1]$ is the contribution of the normal distribution.
@@ 3.4 Summary of the results
Some ensemble properties of spin noise (SN) are similar to DM noise, whereas some properties are different. For both SN and DM, for all models, the area of the parameterised prior is shrunk as much as possible to achieve the maximum posterior probability (given a fixed area, the narrower the distribution is - the higher it’s maximum), but it also has to be sufficiently wide to be consistent with all the samples of noise amplitudes and spectral indices. For spin noise, the normal distribution fits the data as well as the uniform distribution. However, the covariance between $\lg A_{\text{SN}}$ and $\gamma_{\text{SN}}$ plays an important role. In Figures 1(a) and 2(a), one may notice a diagonal trend which results in finding a covariance parameter $\rho_{\text{SN}}\approx-1$. The model with a lack of covariance, $\rho_{\text{SN}}=0$ is disfavored with $\ln\mathcal{B}=6$. The covariance between $\lg A_{\text{SN}}$ and $\gamma_{\text{SN}}$ can be explained by noticing that it follows a line of equal noise power. In contrast, for DM noise, we do not find strong evidence for covariance between $\lg A_{\text{DM}}$ and $\gamma_{\text{DM}}$. With a lack of covariance, the uniform model is strongly preferred over the Gaussian model with a log Bayes factor of $6$. The mixture model is disfavored by our observations of both spin noise and DM noise, indicating the lack of outlying noise terms.
Given the posterior biases found in simulated data studies in Appendix A, we repeated all analyses from this section simultaneously with modelling the CP amplitude $\lg A_{\text{CP}}$ to be drawn from the truncated Normal distribution with hyperparameters $(\mu_{\lg A_{\text{CP}}},\sigma_{\lg A_{\text{CP}}})$. This is the approach that mitigated posterior biases in the simulations. SN and DM noise hyperparameters inferred this way are consistent with those in Table 2 at $1\sigma$ level, except for the mixture model’s $\mu_{\gamma_{\text{SN}}}=1.44^{+0.62}_{-0.84}$. Because the mixture model is disfavored by the data and the rest of the results are consistent with Table 2, we do not report the results obtained under this alternative model for other hyperparameters.
Table 2: Results of hyperparameter estimation and hierarchical model selection for ensemble pulsar noise properties of the European Pulsar Timing Array. DM noise and spin noise are parameterised by $\bm{\theta}=(\lg A,\gamma)$. We present the results for the following models of pulsar noise parameter distributions: a uniform distribution $\mathcal{U}$, a normal distribution $\mathcal{N}$ with and without covariance between $\lg A$ and $\gamma$, a mixture model of a normal and a uniform distribution. Columns $\mathcal{P}(\bm{\Lambda}|\bm{\delta t})$ contain maximum-aposteriori hyperparameter values with $1\sigma$ credible levels. Columns $\ln\mathcal{B}$ contain log Bayes factors in favour of a model of interest against the uniform distribution model. The fact that they are negative means that the uniform model is the best fit.
@@
$\sigma_{\gamma_{\text{SN}}}$ $1.48^{+0.86}_{-0.44}$ $\sigma_{\gamma_{\text{DM}}}$ $0.97^{+0.40}_{-0.24}$
@@ 4 Caveats of the circular analysis
Circular analysis or “double dipping” refers to the usage of data to inform on the model which is then applied to analyse the same data. The methodology we propose in our study is not a circular analysis, it represents one global fit to PTA data. Despite the importance sampling involves performing parameter estimation on the same data twice, the final result is ultimately independent of the proposal distribution obtained at the first step. Namely, the target distribution is independent of the proposal distribution. However, because circular analysis has been extensively discussed in the PTA community, in this section we take an opportunity to cover this subject from the hierarchical inference standpoint.
Overall, circular analyses may lead to the underestimation of measurement uncertainty and systematic errors. However, some PTA analyses in the past have involved a circular analysis, as pointed out by van Haasteren (2025). Namely, pulsar spin noise and DM noise terms were included (not included) in the full-PTA analysis based on (a lack of) evidence for these terms in single-pulsar noise model selection. van Haasteren (2025) argues against this approach. Indeed, double dipping is not a proper statistical treatment, it should be abandoned in favour of the proper model selection and model averaging performed in one global fit, as recommended by the author. However, to give a fair overview of the problem, we also show that circular analysis does not always yield incorrect conclusions in analyses of PTA data.
Before we discuss the application of hierarchical inference in the subsections below, we would like to point out that it is not clear from van Haasteren (2025) that the aforementioned approach of selecting spin noise terms for subsequent full-PTA analyses has led to underestimation of measurement uncertainties reported by PTAs. For model averaging, the prior odds between the existence or absence of CP are proposed to be 50%-50% in van Haasteren (2025). For the case of including a red noise term to every pulsar, the prior odds will be arbitrary and vary from pulsar to pulsar which is clear from the equations in the Appendix in van Haasteren (2025). However, data informs on these odds, so they should become a model (hyper-)parameter to avoid prior misspecification, the same problem discussed in this work. These parameterised odds will act as model selection. In the limit where there is no measurable red noise in pulsars and the data informs on it sufficiently well, this will lead to the elimination of the contribution of one of the models. Because different noise models lead to different measurement uncertainties for $(\lg A_{\text{CP}},\gamma_{\text{CP}})$, it is possible that the uncertainty introduced by enforcing red noise to every pulsar is suboptimal.
@@ 4.1 A comparison between circular analysis, incorrect priors, and a proper analysis
One example of circular analysis in the context of our hierarchical inference is finding best-fit $\bm{\Lambda}$ and using it for the gravitational wave search. It also leads to resolving prior misspecification but potentially at the cost of a reduced measurement uncertainty for $(\lg A,\gamma)$ of the gravitational wave background. The reduction may come thanks to simply ignoring a part of the intrinsic measurement uncertainty for $\bm{\Lambda}$, so a regular noise fluctuation is more likely to render our estimate of $(\lg A,\gamma)$ to be inconsistent with the true value. In Figure 3, we show three measurements of $(\lg A,\gamma)$ of the CP in the EPTA data. Blue dashed contours correspond to a posterior obtained with the standard uniform spin noise priors, $\pi(\bm{\theta}|\bm{\Lambda})=\mathcal{U}(\bm{\theta}|\varnothing)$. Red dotted contours correspond to $\pi(\bm{\theta}|\bm{\Lambda})=\mathcal{N}(\bm{\theta}|\bm{\mu}_{0},\bm{\sigma}_{0},\rho=0)$, where $(\bm{\mu}_{0},\bm{\sigma}_{0})$ are the values obtained for this model from Table 2. Therefore, the red dotted line corresponds to a circular analysis. Filled green contours correspond to a proper (global-fit) measurement obtained based on Equation 9, it is discussed in more detail in Goncharov et al. (2024). The measurement uncertainty of $(\lg A,\gamma)$ obtained with circular analysis matches that obtained with a full hierarchical analysis. It is even $5$-$8$% larger, suggesting that in the current setting circular analysis overestimates the intrinsic measurement uncertainty. Note, however, that the circular analysis shifts the maximum-aposteriori value of $(\lg A,\gamma)$, so the use of it is still not recommended.
@@
Figure 3: A comparison between circular analysis (red dotted contours), incorrect broad uniform priors (blue dashed contours), and a proper hierarchical analysis (green contours) of the common-spectrum process (CP) associated with the gravitational wave background (GWB).
4.2 A toy example: circular analysis of CP
@@
Keeping in mind the caveats above, for the test purposes we also perform the estimation of hyperparameters $\bm{\Lambda}$ for the mixture model of two Gaussian distributions where hyper-priors on the second component of the mixture correspond to EPTA measurement of the CP parameters EPTA Collaboration & InPTA Collaboration (2023b). This exercise remains useful to explore the interplay between pulsar-intrinsic noise and the CP and to understand the performance of the model. For the two-Gaussian mixture model, $\pi(\lg A_{\text{SN}},\gamma_{\text{SN}}|\bm{\Lambda}_{2\mathcal{N}})$, and $\bm{\Lambda}_{2\mathcal{N}}=(\mu_{\lg A_{\text{SN}}}^{1,2}$, $\sigma_{\lg A_{\text{SN}}}^{1,2}$, $\mu_{\gamma_{\text{SN}}}^{1,2}$, $\sigma_{\gamma_{\text{SN}}}^{1,2}$, $\rho_{\text{SN}}^{1,2}$, and $\nu_{\mathcal{N}})$. Indices $(1,2)$ refer to each of the two components in a mixture, while $\nu$ corresponds to a fraction of the first mixture component in the total prior probability. For the second component to correspond to the CP in the EPTA data, we impose a Gaussian hyperprior determined by the measurement uncertainty on $(\lg A,\gamma)$ reported by the EPTA for our data. We find that the fraction of the second normal component is mostly consistent with one and inconsistent with zero. However, hyperparameters of the second mixture component – mean and (co-)variance – are well-constrained. Interestingly, we also find a marginal excess posterior density at $(\sigma_{\lg A_{\text{SN}}}^{2},\sigma_{\gamma_{\text{SN}}}^{2})=(0,0)$. This result is related to Goncharov et al. (2022)’s measurement of $\sigma_{\lg A}$ to be consistent with zero. When removing the CP term with $\gamma=13/3$ from our pulsar-intrinsic noise model, the posterior density increases, as expected. This is shown in Figure 4.
@@
Figure 4: Modelling of the common-spectrum process (CP) as a narrow normally-distributed cluster of pulsar noise using data-informed hyperpriors $\pi(\bm{\Lambda})$. We assume that pulsar noise parameters $\lg A$ and $\gamma$ are distributed as a mixture model of two Gaussians, where hyperparameters of the second Gaussian are chosen to match the measurement uncertainty on the common spectrum process reported by the EPTA Collaboration & InPTA Collaboration (2023b). Parameters $\sigma$ with respective indices correspond to standard deviations of the distribution that governs $\lg A$ and $\gamma$. Consistency of $\sigma$ with zero has first been shown in Goncharov et al. (2022) using a more rigorous although less flexible methodology. Parameter $\nu_{\mathcal{N}}$ corresponds to the fraction of the first, wider normal distribution in a model. Solid lines correspond to including the second $\gamma=13/3$ red noise term to pulsars to account for the presence of the CP, dashed lines correspond to not including it. As expected, the contribution of the CP becomes more visible in the latter case at $\sigma_{\lg A,\gamma}=0$ and $\nu_{\mathcal{N}}\approx 0.3$.
@@ 5 Conclusions
We performed inference of ensemble properties of spin noise and DM noise in the 25-year version of the second data release of the European Pulsar Timing Array (EPTA). Overall, we find that the standard uniform priors used in the previous analysis are too wide and not representative of the observations. However, a parameterised uniform prior distribution is a good fit. We, therefore, recommend correctly accounting for ensemble pulsar noise properties using Equation 9 in future full-PTA analyses to avoid systematic errors (parameter estimation and model selection biases). Our approach of prior reweighting for inferring noise hyperparameters based on Equation 8 is well-suited for understanding broad ensemble noise properties for simulating data and predicting the PTA noise budget.
@@ Acknowledgements
The code to reproduce the results of this analysis and the analysis by Goncharov et al. (2022) is available at github.com/bvgoncharov/pta_priors. The data used for our study, EPTA DR2 by the (EPTA Collaboration, 2023), is available at zenodo. Other data may be provided by the corresponding author upon request. The PTA likelihood is incorporated in enterprise (Ellis et al., 2020) and posterior sampling is performed using ptmcmcsampler (Ellis & van Haasteren, 2017) and dynesty (Speagle, 2020).
Appendix A Tests with the simulated data
We test our inference of hyperparameters based on Section 2.2.1 using two simulated data sets. Both simulations are based on $25$ pulsars, timed over $15$ years with the $100$ ns precision. The cadence is empirically chosen to visually represent irregular PTA observations. It is based on the Pareto distribution with the shape value of $0.6$, scaled to the PTA observation time, with the maximum observation gap of $0.6$ of the time span. Simulations contain spin noise and the common-spectrum process (CP). Spin noise is drawn from the truncated Normal distribution, with hyperparameters given by the first five columns in Table 3. CP represents an additional red noise term in every pulsar, with the same amplitude and spectral index listed in the last two columns in Table 3. Therefore, CP shows the expected spectral properties of the GWB from supermassive black hole binaries.
Table 3: Parameters of our two simulations.
$\mu_{\lg A}$ $\sigma_{\lg A}$ $\mu_{\gamma}$ $\sigma_{\gamma}$ $\rho$ $\lg A_{\text{CP}}$ $\lg\gamma_{\text{CP}}$
$-14.5$ $1.5$ $4.0$ $1.5$ $0.0$ $-15.0$ $13/3$
$-13.8$ $0.7$ $2.0$ $1.0$
The simulations are designed to test our ability to correctly model the distribution of spin noise parameters given the presence of CP. In the first simulated data set, hyperparameters are chosen such that spin noise amplitudes and spectral indices significantly overlap with those of the CP. Because CP and spin noise are indistinguishable in the data of single pulsars, the proposal posteriors we construct from such data may show significant differences from the target posteriors we would get in a global fit to all pulsar data. As described in Section 2.2.1, we fix the spectral index of CP to $13/3$. Thus, in pulsars where $\gamma_{\text{SN}}$ is close to $13/3$, proposal posteriors on $A_{\text{CP}}$ peak at values of $A_{\text{SN}}$, with only the tails extending to the true value of $A_{\text{CP}}$. These biases are mitigated when obtaining target posteriors in the global fit, but the residual biases may remain. In the second simulated data, the overlap between CP and spin noise is less significant. Therefore, the first simulation is considered to be the most difficult case.
First, we perform (hyper-)parameter estimation assuming no covariance between spin noise amplitudes and spectral indices, as simulated ($\rho=0$). The results for the first simulation are shown in Figure 5, and for the second simulation in Figure 6, both as orange contours. The inferred hyperparameters are fully consistent with the simulated values, except for the standard deviation of the spin noise amplitudes $\sigma_{\lg A}$ for the first simulation. The simulated value of $1.5$ lies at the edge of the $1\sigma$ credible level in two-dimensional posteriors including $\sigma_{\lg A}$, and just outside of the $1\sigma$ level in the one-dimensional posterior for $\sigma_{\lg A}$.
Second, we infer hyperparameters assuming there may be a covariance between spin noise amplitudes and spectral indices ($\rho$ is a free parameter). The results are shown as green contours in Figure 5 and Figure 6. Although we simulated $\rho=0$, for both simulations we find $\rho$ to be inconsistent with 00. Instead, the maximum-aposteriori $\rho$ is consistent with $-1$, while the value of 00 is beyond $3\sigma$ credibility. Compared to the previous case where we fix $\rho=0$, a minor tension at $1$-$2\sigma$ level emerges for $\mu_{\gamma}$ in the second simulation, but another minor tension for $\sigma_{\lg A}$ in the first simulation disappears. We conclude that $\rho$ is most sensitive to posterior biases.
To resolve posterior biases, we extend our model to $A_{\text{CP}}$, imposing that it is also drawn from the truncated Gaussian distribution. Thus, we extend both spin noise and CP priors to become a part of our model: $\pi(\lg A,\gamma|\mu_{\lg A},\sigma_{\lg A},\mu_{\gamma},\sigma_{\gamma},\rho)\pi(\lg A_{\text{CP}}|\mu_{\lg A_{\text{CP}}},\sigma_{\lg A_{\text{CP}}})$. We impose a Log-Uniform prior on $\sigma$ parameters. The results are shown as black contours in Figure 5 and Figure 6. In this case, we find the inferred hyperparameters to be fully consistent with the simulated values. The values of $\mu_{\lg A_{\text{CP}}}$ and $\sigma_{\lg A_{\text{CP}}}$ are also in accordance with our expectations. The value $\mu_{\lg A_{\text{CP}}}$ is consistent with the simulated $\lg A_{\text{CP}}=-15.0$. The value of $\sigma_{\lg A_{\text{CP}}}$ peaks at small values, it excludes 00 in the first simulation, despite CP representing an infinitely small range of amplitudes. This is a known effect in importance sampling due to a finite number of the recycled posterior samples (Goncharov et al., 2022).
Figure 5: Posteriors on the inferred hyperparameters in the first simulation. Dashed vertical lines correspond to simulated values in Table 3. Yellow contours corresponds to $\rho=0$, green contours correspond to $\rho$ as a free parameter, and black contours correspond to additionally modelling the distribution of CP amplitudes in pulsars as being drawn from the truncated Normal distribution with mean $\mu_{\lg A_{\text{CP}}}$ and the standard deviation $\sigma_{\lg A_{\text{CP}}}$. Shaded areas correspond to $1\sigma$ ($1\sigma$ and $2\sigma$) credible levels in fully-marginalised posteriors (two-dimensional posteriors).
Figure 6: Posteriors on the inferred hyperparameters in the second simulation. Dashed vertical lines correspond to simulated values in Table 3. Colours and contours are explained in Figure 5.
Appendix B Numerical marginalisation over hyperparameters